OpenStudy (anonymous):

Will give medal and fan Please help Find the sum and product of the roots. 3x^2 - 4x - 7 = 0 A: sum= 4/3, product= 7/3 B: sum= -4/7, product= -7/3 C: sum= -4, product= -7 D: sum= 3/5, product= 7/3

OpenStudy (anonymous):

for \[ax^2+bx+c=0\] the sum of the roots is \(-\frac{b}{a}\) and the product of the roots is \(\frac{c}{a}\)

OpenStudy (anonymous):

so that would be C? correct

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

\[-\frac{b}{a}=-\frac{-4}{3}=\frac{4}{3}\]

OpenStudy (anonymous):

oh i ment B but that was still wrong haha its A?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

<3